# Learning Force Distribution Estimation for the GelSight Mini Optical Tactile Sensor Based on Finite Element Analysis

### v2

@@ I Introduction
Tactile sensing plays an important role in advancing the state-of-the-art in robotic manipulation [1, 2, 3, 4, 5, 6, 7, 8]. Successful applications include grip adaptation through slip detection [9, 10, 11, 12], medical procedures [13, 14] and tele-operation [15].
In particular, optical tactile sensors have emerged as a promising technology for capturing contact information due to their high spatial resolution, multimodal sensing capabilities [16]—including shape [17], hardness [18], texture [19], and temperature [20]—and cost-effectiveness [21, 14]. However, many prior works have focused on extracting only low-dimensional tactile information, such as total force [22, 23, 24], limiting operational flexibility. Access to contact force distributions, on the other hand, would enable better handling of multiple contacts and diverse manipulation scenarios [25].
@@
To summarize, our key contributions are i) a method for creating force labels from FEA outputs tailored to the commercially available GelSight Mini, ii) the collection of a large dataset of versatile indentation data, iii) a model architecture capable of generalizing across objects and allowing for real-time inference, and iv) open-sourcing our dataset, the code for data collection, label generation, and model training, as well as the trained models for aiding reuse and reproducibility.
@@ II-A Marker Displacement Methods
Li et al. [38] posit that it is the contact layer deformations that capture the crucial information within tactile images. By analyzing these deformations, various contact features can be extracted, with Marker Displacement Methods being the most common approach [38]. In MDMs, markers are placed on or within the elastomer and appear as features in the sensor’s imagery (Fig. 1). For the GelSight sensor [30, 35], markers were first introduced in [40] to study normal and shear forces, along with slip dynamics. They identified a linear relationship between loads and marker motion, but this applied only in non-slip conditions. Beyond marker motion, optical sensors can capture detailed height maps and contact geometry through careful illumination and photometric stereo [39]. These height maps can be used to estimate contact forces with a third-degree polynomial [15].
In this paper, we use a gel with markers, but their movement is not explicitly tracked. Instead, they serve as implicit features within the sensor image, which is analyzed by a neural network to predict force distributions.
II-B Deep Learning-Based Tactile Representations
Advancements in computer vision directly translate to vision-based tactile sensing. Models such as CNNs and LSTMs were adapted to assess object hardness [18] and grip stability [41], whereas SVMs were used for lump detection [42]. More tactile-specific deep learning methods have been developed for overall force prediction [30] and for pixel-wise contact area and normal force estimation [31].
@@ II-C Force Distribution Prediction Through Elasticity Theory
However, direct prediction of force distributions from displacements, usually through a linear stiffness matrix, does not account for the nonlinearities of soft elastomers. Sun et al. [44] addressed this limitation by employing ResNet [45], which was trained on sensor images with approximated force distributions. Similarly, Sferrazza et al. [33] utilized a Deep Neural Network (DNN) trained on image features with force distributions obtained from FEA, and later improved this method by incorporating simulated training data [25, 32]. A similar approach was also adopted in [34], where the authors additionally investigated soft objects interacting with the tactile sensor.
Building on these approaches, we also aim to estimate contact forces using supervised deep learning. However, instead of working on optical flow features or grayscale images [32] applicable to the custom-made sensor with a dense 3D-marker field [25, 32, 33], we use raw RGB images from a widely available GelSight Mini sensor. Crucially, we develop specific procedures for data collection and model training that enable the efficient use of this widely accessible sensor, thereby significantly lowering the entry barrier into the field.
Our key contributions are: i) method for creating force labels from FEA outputs tailored to GelSight Mini + implementation in CalculiX, ii) data collection procedure + dataset, iii) trained model applicable to varying objects and gels.
@@ III-A Data Collection
The data collection process involves a series of precise indentation experiments. We attach a GelSight Mini sensor to the fixed spindle of a Computer Numerical Control (CNC) milling machine, with a positional tolerance of ±$\pm 0.25\,\mathrm{\mu m}$ to automatically create random contact configurations between the sensor and selected indenters (cf. Fig. 2). This work leverages $24$ 3-d printed indenters ($19$ for training and $5$ for testing) varying in shape and size as shown at the bottom of Fig. 2. The optical tactile sensor, i.e., Gelsight Mini, is equipped with the dotted gel to better track the indentation motion. We additionally place a six-axis RESENSE-HEX-21 Force/Torque (F/T) sensor above the Gelsight Mini to have a complementary external force measurement, which is later used to validate the material model. With this setup, we collected a total of $14279$ samples. Each sample contains the respective GelSight Mini RGB image, the CNC motion data, i.e., the current indentation, and the F/T sensor’s readings. The forces in the $z$ direction reach up to $40\,\mathrm{N}$, and in the $x$ and $y$ directions up to ±$\pm 5\,\mathrm{N}$.
@@ III-C Creating Labels
To obtain the “ground truth” force distribution labels, each real-world indentation experiment (cf. Sec. III-A) is repeated in simulation in order to calculate the contact forces using FEA (cf. Sec. III-B & Fig. 3). For each element on the surface of the gel, three force components are computed: shear forces in the $x$ and $y$ directions, and normal forces in the $z$ direction.
@@ III-D U-Net for Learning Force Distribution Estimation
We train the model using $9084$ samples and use a validation dataset containing $538$ samples. We later evaluate the model on three different test datasets for which we provide the details in the respective experimental sections. Prior to training, both input images and the ground truth force distribution labels are normalized with respect to the training dataset to ensure a consistent baseline across the data. Shear forces are normalized to a range of $-1$ to $+1$, while normal forces range from 00 to $+1$. Data augmentation techniques, such as adding Gaussian noise and adjusting image brightness, contrast, saturation, and hue, are used to enhance the model’s generalization capabilities. Training is carried out using the Adam optimizer [47], with an initial learning rate of $0.001$ and a batch size of $8$. The learning rate is adjusted adaptively based on validation performance. As validation loss, the Mean Absolute Error (MAE) on the sum of total forces in the unnormalized space is used. We train for $400$ epochs and select the model with the lowest validation loss.
Figure 4: U-net model which maps raw images from GelSight Mini sensor to shear and normal force distributions. The architecture comprises $4$ down-sampling (encoder) and $4$ up-sampling (decoder) blocks, connected by skip connections. The number of feature channels at each stage is labeled at the bottom of the corresponding block. Different colors of the boxes and arrows indicate specific operations and activation functions. This image is generated using PlotNeuralNet [48].
@@ IV-A Material Characterization
To validate the Neo-Hookean material model parameter $C_{10}$, load-depth indentation data can be utilized (cf. Sec. III-B). The load-depth measurements are obtained, using a sphere indenter with a $15\,\mathrm{mm}$ diameter, by sampling data points at different indentation depths ranging from 0.5 mm to 2.0 mm, increasing in 0.5 mm intervals. The measured normal forces from the F/T sensor in our experimental setup $\tilde{f}^{(i)}_{z}$ are compared with the forces $f^{(i)}_{z}(C_{10})$ calculated by the FEA via the MSE loss
$J(C_{10})=\frac{1}{N}\sum_{i=1}^{N}\Bigl{(}f^{(i)}_{z}(C_{10})-\tilde{f}^{(i)}_{z}\Bigr{)}^{2}.$ (2)
When running Bayesian optimization to find the best fit for these $N=4$ measurements, we found that $\hat{C}_{10}=0.0792$ provided the best fit with a MAE of $0.5166\,\mathrm{N}$ between FEA and F/T sensor measurements. Thus, we confirm that our estimate is within the range of $C_{10}=0.0725$ previously reported in [30] for the GelSight sensor. In the experiments, we use that value as it was stated in the original GelSight paper and our estimate closely matches it.
@@ IV-B Evaluation of U-net Force Distribution Estimation
Next, we assess the U-nets’ capacity to accurately estimate shear and normal force distributions from the raw RGB images of a GelSight Mini sensor. We evaluate the trained networks on multiple test datasets that cover a wide range of forces and contact configurations. The test dataset collection follows the same procedure as described in Secs. III-A & III-C).
@@
Throughout all the experiments, to quantitatively evaluate the performance of predicting force distributions, we employ the MAE metric. With MAE, the deviations between the predicted and ground truth force values can be measured, both for per-Grid Unit Forces (GUF) and for the Total Force (TF) values. We are reporting the GUF errors in addition to the TF errors since they directly reflect the error in the models’ individual output values. They, therefore, also capture spatial information regarding the contact area, whereas TF is obtained as the sum of the GUFs and is therefore missing out on the local information.
@@
[…]
@@
While the previous experiment evaluated our trained representation on the same indenters that were also used for training, this section now evaluates FEATS on $5$ new, previously unseen test indenters which were not included in the training data (cf. Fig. 2). We collected $2789$ samples with these new indenters, covering the same force range as for the other testing dataset (cf. Fig. 6). The results in Tab. III reveal that our trained network is indeed capable of generalizing to the previously unseen indenters. Considering the per-Grid Unit Forces (GUF), the errors only slightly increase by $0.0004\,\mathrm{N}$, $0.0003\,\mathrm{N}$, for the x- and y-direction, while the total force errors remain comparable to the previous experiment. In line with the higher range of the normal forces in this dataset, we also observe a slightly increased error of $0.0012\,\mathrm{N}$ compared to the evaluation on the known indenters. Since we only observe a small increase in the errors compared to the known indenters, we conclude that the representation generalizes well to these new contact configurations.
@@
To explore the generalization capability of our FEATS method, we finally test it on images captured with a different GelSight Mini sensor of the same type. We again consider the previously unseen test indenters. As shown in Fig. 7, qualitatively, the predictions remain largely accurate. Quantitatively, we obtain mean total force errors of $0.24\,\mathrm{N}$, $0.27\,\mathrm{N}$, & $2.99\,\mathrm{N}$, for the $x$-, $y$-, and & $z$-direction, respectively. Similar to the results of the previous sections, we therefore again observe the largest increase in error for the normal force ($z$-direction), where also the data range is substantially larger. We hypothesize that the additional increase in error might be related with several factors, including differences in the alignment of the camera relative to the gel, and sensor-related factors such as image brightness, contrast, saturation and hue.
@@ IV-C Inference Speed
To determine the speed of the U-net, we measure the time it takes to process a single input on a NVIDIA Quadro RTX 5000 GPU and a Intel Core i7-10875H 8-Core CPU with a frequency of $2.3\,\mathrm{GHz}$. The time required for transferring data between the CPU and GPU, as well as asynchronous execution and warm-up of the GPU is taken into consideration. On average, the U-net achieves an inference time of $4.1746\pm 1.339\,\mathrm{ms}$ over $300$ runs, which is sufficient for obtaining the force distributions in real-time since the GelSight Mini sensor runs at $25\,\mathrm{Hz}$. This is significantly faster compared to running the FEA simulations, which may take $10-120$ minutes depending on the contact geometry.
@@ V Conclusion
We introduced FEATS—a machine learning approach for estimating force distributions using the GelSight Mini tactile sensor. By training a U-net model on FEA-derived data, we achieved accurate predictions of both shear and normal forces from raw sensor images, with Mean Absolute Error (MAE) under $0.8\,\mathrm{N}$ on average in the range $0-40\,\mathrm{N}$ for the total normal force and under $0.3\,\mathrm{N}$ on average for the total shear force in each direction with a range of $-5\,\mathrm{N}$ to $5\,\mathrm{N}$. The model shows potential for real-time applications and generalization to different indenters and sensors, although additional improvements are needed to account for the domain gap when a different sensor is used. This could be addressed, e.g., via the recently proposed domain adaptation for optical tactile sensors [49]. Our FEATS method offers an efficient online approach that amortizes the cost of running inverse FEA into a forward pass of a neural network, resulting in physically-grounded, interpretable representations for optical tactile sensors. Future work will focus on enhancing the model’s robustness to different tactile scenarios, refining its generalization capabilities across a broader range of objects and sensor types, and work towards demonstrating the representation’s effectiveness in robotic manipulation tasks.

### v3

@@ I Introduction
Tactile sensing plays an important role in advancing the state-of-the-art in robotic manipulation [1, 2, 3, 4, 5]. Successful applications include grip adaptation through slip detection [6, 7, 8, 9], medical procedures [10, 11] and tele-operation [12]. In particular, optical tactile sensors have emerged as a promising technology for capturing contact information due to their high spatial resolution, multimodal sensing capabilities [13]—including shape [14], hardness [15], texture [16], and temperature [17]—and cost-effectiveness [18, 11]. However, many prior works have focused on extracting only low-dimensional tactile information, such as total force [19, 20, 21], limiting operational flexibility. Access to contact force distributions, on the other hand, would enable better handling of multiple contacts and diverse manipulation scenarios [22].
@@
To summarize, our key contributions are i) a method for creating force labels from FEA outputs tailored to the commercially available GelSight Mini, ii) the collection of a large dataset of versatile indentation data, iii) a model architecture capable of generalizing across objects and allowing for real-time inference, and iv) open-sourcing our dataset, the code for data collection, label generation, model training, and the trained models for aiding reuse and reproducibility.
@@ II-A Marker Displacement Methods
[…]
II-B Learning-Based Tactile Representations
Advancements in computer vision directly translate to vision-based tactile sensing. Models such as CNNs, LSTMs, and SVMs were adapted to assess object hardness [15], grip stability [38], and lump detection [39]. More tactile-specific deep learning methods have been developed for overall force [27] and pixel-wise contact area and normal force estimation [28].
@@ II-C Force Distribution Prediction Through Elasticity Theory
However, direct prediction of force distributions from displacements, usually through a linear stiffness matrix, does not account for the nonlinearities of soft elastomers. Sun et al. [41] addressed this limitation by training a ResNet [42] on sensor images with approximated force distributions using a thumb-sized vision-based tactile sensor called Insight. Andrussow et al. [43] presented a miniaturized successor, Minisight, which retains the same force labeling methodology in a fingertip-sized form factor. In contrast, Sferrazza et al. [30] utilized a neural network trained on image features with force distributions obtained from FEA, providing more physically accurate labels. They subsequently improved this approach using simulated training data [22, 29]. The use of FEA-generated labels was further demonstrated by Xue et al. [31], which also considered interactions with soft objects. This methodology was later adopted for larger tactile surfaces, with Luu et al. [44] applying it to the TacLINK [25] sensor, and Ho and Nakayama [45] incorporating it in their IoTouch system.
Building on these approaches, we also aim to estimate contact forces using supervised deep learning. However, in contrast to the aforementioned approaches relying on pipelines designed for non-public sensors and involving preprocessing (e.g., optical flow [29]) or sim-to-real transfer [44], we use raw RGB images from the commercially available GelSight Mini sensor. Crucially, we develop specific procedures for data collection and model training that enable the efficient use of this widely accessible sensor, thereby significantly lowering the entry barrier into the field.
Our key contributions are: i) method for creating force labels from FEA outputs tailored to GelSight Mini + implementation in CalculiX, ii) data collection procedure + dataset, iii) trained model applicable to varying objects and different sensors of the same type.
@@ III-A Data Collection
The data collection process involves a series of precise indentation experiments. We attach a GelSight Mini sensor to the fixed spindle of a Computer Numerical Control (CNC) milling machine, with a positional tolerance of $\pm 0.25\,\mathrm{\mu m}$ to automatically create random contact configurations between the sensor and selected indenters (cf. Fig. 2). This work leverages $24$ 3-d printed indenters ($19$ for training and $5$ for testing) varying in shape and size as shown at the bottom of Fig. 2. The optical tactile sensor, i.e., Gelsight Mini, is equipped with the dotted gel to better track the indentation motion. We additionally place a six-axis RESENSE-HEX-21 Force/Torque (F/T) sensor above the Gelsight Mini to have a complementary external force measurement, which is later used to validate the material model. With this setup, we collected a total of $14279$ samples. Each sample contains the respective GelSight Mini RGB image, the CNC motion data, i.e., the current indentation, and the F/T sensor’s readings. The forces in the $z$ direction reach up to $40\,\mathrm{N}$, and in the $x$ and $y$ directions up to $\pm 5\,\mathrm{N}$.
@@ III-C Creating Labels
To obtain the approximated “ground truth” force distribution labels, each real-world indentation experiment (cf. Sec. III-A) is repeated in simulation in order to calculate the contact forces using FEA (cf. Sec. III-B & Fig. 3). For each element on the surface of the gel, three force components are computed: shear forces in the $x$ and $y$ directions, and normal forces in the $z$ direction.
@@ III-D U-Net for Learning Force Distribution Estimation
We train the model using $9084$ samples and use a validation dataset containing $538$ samples. We later evaluate the model on three different test datasets for which we provide the details in the respective experimental sections. Prior to training, both input images and the ground truth force distribution labels are normalized with respect to the training dataset to ensure a consistent baseline across the data. Shear forces are normalized to a range of 1-1 to +1+1, while normal forces range from 0 to +1+1. Data augmentation techniques, such as adding Gaussian noise and adjusting image brightness, contrast, saturation, and hue, are used to enhance the model’s generalization capabilities. Training is carried out using the Adam optimizer [47], with an initial learning rate of $0.001$ and a batch size of $8$. The learning rate is adjusted adaptively based on validation performance. As validation loss, the Mean Absolute Error (MAE) on the sum of total forces in the unnormalized space is used. We train for $400$ epochs and select the model with the lowest validation loss.
Figure 4: U-net model which maps raw images from GelSight Mini sensor to shear and normal force distributions. The architecture comprises 44 down-sampling (encoder) and 44 up-sampling (decoder) blocks, connected by skip connections. The number of feature channels at each stage is labeled at the bottom of the corresponding block. Different colors of the boxes and arrows indicate specific operations and activation functions. This image is generated using PlotNeuralNet [48].
@@ IV-A Material Characterization
To validate the Neo-Hookean material model parameter $C_{10}$, load-depth indentation data can be utilized (cf. Sec. III-B). The load-depth measurements are obtained, using a sphere indenter with a $15\,\mathrm{mm}$ diameter, by sampling data points at different indentation depths ranging from 0.5 mm to 2.0 mm, increasing in 0.5 mm intervals. The measured normal forces from the F/T sensor in our experimental setup $\tilde{f}^{(i)}_{z}$ are compared with the forces $f^{(i)}_{z}(C_{10})$ calculated by the FEA via the MSE loss $J(C_{10})=1/N\sum_{i=1}^{N}\left(f^{(i)}_{z}(C_{10})-\tilde{f}^{(i)}_{z}\right)^{2}$. When running Bayesian optimization to find the best fit for these $N=4$ measurements, we found that $\hat{C}_{10}=0.0792$ provided the best fit with a MAE of $0.5166\,\mathrm{N}$ between FEA and F/T sensor measurements. Thus, we confirm that our estimate is within the range of $C_{10}=0.0725$ previously reported in [27] for the GelSight sensor. In the experiments, we use that value as it was stated in the original GelSight paper and our estimate closely matches it.
@@ IV-B Evaluation of U-net Force Distribution Estimation
Next, we assess the U-nets’ capacity to accurately estimate shear and normal force distributions from the raw RGB images of a GelSight Mini sensor. We evaluate the trained networks on multiple test datasets that cover a wide range of forces and contact configurations. The test dataset collection follows the procedure described in Secs. III-A & III-C).
@@
Throughout all the experiments, we quantitatively evaluate the model performance through reporting the MAE between the predicted and ground truth force values, both for per-Grid Unit Forces (GUF) and for the Total Force (TF) values, i.e.,
$\text{MAE}_{\text{GUF}}(\mathbf{f},\mathbf{\hat{f}})=\frac{1}{3\text{WH}}\sum_{i=0}^{\text{W}}\sum_{j=0}^{\text{H}}\left|\mathbf{f}^{(i,j)}-\hat{\mathbf{f}}^{(i,j)}\right|$
$\text{MAE}_{\text{TF}}(\mathbf{f},\mathbf{\hat{f}})=\left|\sum_{i=0}^{\text{W}}\sum_{j=0}^{\text{H}}\mathbf{f}^{(i,j)}-\sum_{i=0}^{\text{W}}\sum_{j=0}^{\text{H}}\hat{\mathbf{f}}^{(i,j)}\right|\>.$
We are reporting the GUF errors in addition to the TF errors since they directly reflect the error in the models’ individual output values. They, therefore, also capture spatial information regarding the contact area, whereas TF is missing out on the local information.
@@
$\text{U-net}^{\text{\tiny$12{\times}16{\times}3$}}$ $0.206\pm 0.307$ $0.163\pm 0.242$ $0.761\pm 0.941$
$3{\times}\text{U-net}^{\text{\tiny$24{\times}32{\times}1$}}$ $0.426\pm 0.521$ $0.401\pm 0.572$ $\color[rgb]{1,.5,0}0.783\pm 1.022$}
$\textbf{\text{U-net}}^{\text{\tiny$24{\times}32{\times}3$}}$ (ours) $\color[rgb]{.5,0,.5}0.293\pm 0.417$} $\color[rgb]{.5,0,.5}0.230\pm 0.341$} $\color[rgb]{.5,0,.5}0.799\pm 1.014$}
$\text{U-net}^{\text{\tiny$48{\times}64{\times}3$}}$ $0.338\pm 0.484$ $0.272\pm 0.392$ $1.037\pm 1.287$
@@
While the previous experiment evaluated our trained representation on the same indenters that were also used for training, this section now evaluates FEATS on $5$ new, previously unseen test indenters which were not included in the training data (cf. Fig. 2). We collected $2789$ samples with these new indenters, covering the same force range as for the other testing dataset (cf. Fig. 6). The results in Tab. III reveal that our trained network is indeed capable of generalizing to the previously unseen indenters. Considering the per-Grid Unit Forces (GUF), the errors only slightly increase by $004\,\mathrm{N}, 0.000$$003\,\mathrm{N}$, for the x- and y-direction, while the total force errors remain comparable to the previous experiment. In line with the higher range of the normal forces in this dataset, we also observe a slightly increased error of $0.0012\,\mathrm{N}$ compared to the evaluation on the known indenters. Since we only observe a small increase in the errors compared to the known indenters, we conclude that the representation generalizes well to these new contact configurations.
@@
To explore the generalization capability of our FEATS method, we finally test it on images captured with a different GelSight Mini sensor of the same type and a new gel. We again consider the previously unseen test indenters. Initially, we obtain mean total force errors of $0.24\,\mathrm{N}$, $0.27\,\mathrm{N}$, & $2.99\,\mathrm{N}$ for the $x$-, $y$-, & $z$-directions, respectively. To counteract the increased errors compared to the previous evaluation, we collected $4400$ additional samples from four GelSight Minis, systematically swapping their gels, and integrated these samples into our training dataset. The additional data is intended to capture the variations among different GelSight Mini sensors. This significantly enhanced generalization, achieving reduced errors of $0.23\,\mathrm{N}$, $0.25\,\mathrm{N}$, & $1.79\,\mathrm{N}$ for a previously unseen sensor-gel combination (see Fig. 7). The larger increase in error for the normal force ($z$-direction) might be attributed to factors like camera alignment and sensor-specific image variations. A demonstration video showing a dynamic interaction is available on our project website.
@@ IV-C Inference Speed
We finally report our method’s inference time on a computer equipped with a NVIDIA Quadro RTX 5000 and an Intel Core i7-10875H 8-Core CPU. The time for transferring data between the CPU and GPU, as well as asynchronous execution and warm-up of the GPU is taken into consideration. On average, the U-net achieves an inference time of $4.1746\pm 1.339\,\mathrm{ms}$ over $300$ runs, which is sufficient for obtaining the force distributions in real-time since the GelSight Mini sensor runs at $25\,\mathrm{Hz}$. This is significantly faster compared to running the FEA simulations, which may take 1012010-120 minutes depending on the contact geometry.
@@ V Conclusion
We introduced FEATS—a machine learning approach for estimating force distributions using the GelSight Mini tactile sensor. By training a U-net model on FEA-derived data, we achieved accurate predictions of both shear and normal forces from raw sensor images, with Mean Absolute Error (MAE) under $0.8\,\mathrm{N}$ on average in the range $0-40\,\mathrm{N}$ for the total normal force and under $0.3\,\mathrm{N}$ on average for the total shear force in each direction with a range of $5\,\mathrm{N}$ to $5\,\mathrm{N}$. The model showed potential for real-time applications and generalization to previously unseen indenters and GelSight Mini sensors, which could be further improved through domain adaptation [49]. Our FEATS method offers an efficient online approach that amortizes the cost of running inverse FEA into a forward pass of a neural network, resulting in physically-grounded, interpretable representations for optical tactile sensors. Future work will focus on enhancing the model’s robustness to different tactile scenarios and work towards demonstrating the representation’s effectiveness in robotic manipulation tasks.